# HINT: Hierarchical Inter-frame Correlation for One-Shot Point Cloud Sequence Compression

### v1

@@ Abstract
Deep learning has demonstrated strong capability in compressing point clouds. Within this area, entropy modeling for lossless compression is widely investigated. However, most methods rely solely on parent/sibling contexts and level-wise autoregression, which suffers from decoding latency on the order of $10^{1}-10^{2}$ seconds. We propose HINT, a method that integrates temporal and spatial correlation for sequential point cloud compression. Specifically, it first uses a two-stage temporal feature extraction: (i) a parent-level existence map and (ii) a child-level neighborhood lookup in the previous frame. These cues are fused with the spatial features via element-wise addition and encoded with a group-wise strategy. Experimental results show that HINT achieves encoding and decoding time at 105 ms and 140 ms, respectively, equivalent to 49.6 $\times$ and 21.6 $\times$ acceleration in comparison with G-PCC, while achieving up to bit rate reduction of 43.6%, in addition, consistently outperforming the strong spatial-only baseline (RENO).
@@ 1 Introduction
Fig. 2: Overview of HINT. Two consecutive frames are processed by the Temporal Module to produce a parent-level feature $T_{d}$ and a child-level feature $T_{d+1}$. In parallel, the RENO path provides spatial feature $F_{s}$ from frame $t$ (Prior ResNet at parents). The parent-level features are fused by an element-wise addition and broadcast to child-level voxels. Then, it is fused with $T_{d+1}$ as the final feature. Each parent’s eight children are split by parity into two groups, we first predict $G_{e}$’s probability with the final feature. Ground truth $G_{e}$ is embedded and aggregated with final feature as additional information to predict $G_{o}$’s probability. The predicted probabilities are passed to an entropy encoder to produce the bitstream.
@@ 3 HINT
HINT consists of three primary steps: (i) quantization and hierarchy construction, (ii) temporal and spatial feature extraction, and (iii) entropy coding using estimated probability.
@@ 3.2 Parent-level temporal correlation
where elements in $\mathbf{z}$ indicates the existence at $C_{d}^{(t)}+\delta$ (11 if present, else 0). This operation is computed by one neighborhood expansion and two binary searches over sorted Morton keys of $C_{d}^{(t)}$ and $C_{d}^{(t-1)}$. The “existence map” $M$ is produced by concatenation as follows:
@@ 3.3 Child-level temporal correlation
We define a query offset set $\mathcal{N}_{d+1}$ of size $V_{d+1}$, (default $V_{d+1}=125$, i.e., a window of $5\times 5\times 5$). For each child coordinate $C^{(t)}_{d+1}\!\in\!\mathcal{C}_{d+1}^{(t)}$, we traverse all $\delta\in\mathcal{N}_{d+1}$ and look up $C^{(t)}_{d+1}+\delta$ in the previous-frame set $\mathcal{C}_{d+1}^{(t-1)}$. If found, we read its 8-bit occupancy code $O^{(t-1)}_{d+1}(C_{d+1}^{t}+\delta)\in\{0,\ldots,255\}$, otherwise 0 for absent. Let $E\in\mathbb{R}^{256\times 32}$ be a 256-entry embedding table. For each offset, we take the embedding of occupancy code $O^{(t-1)}_{d+1}$ at $C^{(t)}_{d+1}+\delta$ by $e_{\delta}=E\!\big(O^{(t-1)}_{d+1}(C^{(t)}_{d+1}+\delta)\big)$. We average the neighbor descriptors over $\mathcal{N}_{d+1}$, and apply a linear projection $W_{t}$ to obtain a 32-dim feature:
@@
Table 1: Mean BPP (bits per point) over 300 frames (vox10).
@@
Table 2: Computational complexity comparison on Loot & Redandblack (mean per frame).
@@ 3.4 Entropy coding based on sibling correlation
Before encoding $G_{o}$, we summarize the already decoded even siblings into a lightweight context vector. Let $O_{e}\in\{0,\ldots,255\}$ be the decoded 8-bit code of occupancy code. We embed each code with a 256-entry table $E\in\mathbb{R}^{256\times 32}$ and concatenate with its 3-dim relative position $\pi(i)\in\mathbb{R}^{3}$ (the child’s $(b_{x},b_{y},b_{z})$ inside the parent), then we apply a linear projection $W_{s}\in\mathbb{R}^{32\times(32+3)}$. The four descriptors are averaged with a mask (over available siblings) to form a sibling feature of the same width as $F_{d+1}$:
$F(G_{e})=\frac{1}{|G_{e}|}\sum_{i\in G_{e}}W_{s}\,[E(O_{e}),\,\pi(i)],$ (9)
where $E(O_{e})$ is the embedded occupancy code. For any child voxel $j\in G_{o}$, we fuse this context by element-wise addition for richer context $\tilde{F}_{d+1}$: $\tilde{F}_{d+1}=F_{d+1}+F(G_{e}).$ Then we reuse the same two prediction heads: first predict $s_{0}(j)$ from $\tilde{F}_{d+1}$, then embed $s_{0}(j)$, add it to $\tilde{F}_{d+1}$, and predict $s_{1}(j)$. Arithmetic coding is applied after each prediction. This keeps strict causality: odd children depend only on already decoded even siblings within the same parent.
@@ 4 Experiments
All experiments are run on a desktop with an Intel Core i9-14900 CPU and an NVIDIA RTX-4090 GPU. We implement the codec in PyTorch [14] with TorchSparse [15] for sparse 3D operations. We use the torchac [16] repository for entropy encoding without further optimization.
@@
Main results Tab. 1 summarizes mean BPP at 10-bit precision on 8iVFBv2. HINT is competitive across all sequences: it achieves comparable compression ratio to ECM-OPCC and consistently outperforms G-PCC, SparsePCGC, OctAttention, ACNP-OctAttention, and VoxelDNN. Although with a slight rate gap to ECM-OPCC, Tab. 2 shows that HINT reduces encoding and decoding latency by one to two orders of magnitude compared with ECM-OPCC. This combination of compression and computational efficiency highlights HINT as a codec for practical deployment. Moreover, computational complexity and model size are measured in Tab. 2. While running at 105 ms and 140 ms for encoding and decoding (a $36.2\times$ and $29.3\times$ encoding and decoding speedups over SparsePCGC and $49.6\times$ and $21.6\times$ over G-PCC (encode/decode)), it remains a 3.4 MB (vs. 4.9 MB for SparsePCGC) memory footprint.

### v2

@@ Abstract
Deep learning has demonstrated strong capability in compressing point clouds. Within this area, entropy modeling for lossless compression is widely investigated. However, most methods rely solely on parent/sibling contexts and level-wise autoregression, which suffers from decoding latency on the order of $10^{1}-10^{2}$ seconds. We propose HINT, a method that integrates temporal and spatial correlation for sequential point cloud compression. Specifically, it first uses a two-stage temporal feature extraction: (i) a parent-level existence map and (ii) a child-level neighborhood lookup in the previous frame. These cues are fused with the spatial features via element-wise addition and encoded with a group-wise strategy. Experimental results show that HINT achieves encoding and decoding time at 105 ms and 140 ms, respectively, equivalent to 49.6 $\times$ and 21.6 $\times$ acceleration in comparison with G-PCC, while achieving up to 43.6% bitrate reduction and consistently outperforming the spatial-only baseline (RENO).
@@ 1 Introduction
Fig. 2: Overview of HINT. Frame $t$ and t1t-1 are processed by the Temporal Module to produce a parent-level feature $T_{d}$ and a child-level feature $T_{d+1}$. The RENO path provides spatial feature $F_{s}$ from frame $t$ (Prior ResNet at parents). The parent-level features are fused by an element-wise addition and broadcast to child level. Then, it is fused with $T_{d+1}$ as the final feature. Each parent’s eight children are split by parity into two groups, and we first predict $G_{e}$’s probability with the final feature. Ground truth $G_{e}$ is embedded and aggregated with final feature as additional information to predict $G_{o}$’s probability. The predicted probabilities are passed to an entropy encoder to produce the bitstream.
@@ 3 HINT
HINT consists of three primary steps: (i) optional quantization and hierarchy construction  [^1] A standard approach to process the raw point cloud, but for MPEG 8i Voxelized Full Bodies v2 (8iVFBv2) we skip this step as it has been quantized already. , (ii) temporal and spatial feature extraction, and (iii) entropy coding using estimated probability.
@@ 3.2 Parent-level temporal correlation
where elements in $\mathbf{z}$ indicate the existence at $C_{d}^{(t)}+\delta$ (11 if present, else 0). This operation is computed by one neighborhood expansion and two binary searches over sorted Morton keys of $C_{d}^{(t)}$ and $C_{d}^{(t-1)}$. The “existence map” $M$ is produced by concatenation as follows:
@@ 3.3 Child-level temporal correlation
We define a query offset set $\mathcal{N}_{d+1}$ of size $V_{d+1}$ (default $V_{d+1}=125$, i.e., a window of $5\times 5\times 5$). For each child coordinate $C^{(t)}_{d+1}\!\in\!\mathcal{C}_{d+1}^{(t)}$, we traverse all $\delta\in\mathcal{N}_{d+1}$ and look up $C^{(t)}_{d+1}+\delta$ in the previous-frame set $\mathcal{C}_{d+1}^{(t-1)}$. If found, we read its 8-bit occupancy code $O^{(t-1)}_{d+1}(C_{d+1}^{(t)}+\delta)\in\{0,\ldots,255\}$, otherwise 0 for absent. Let $E\in\mathbb{R}^{256\times 32}$ be a 256-entry embedding table. For each offset, we take the embedding of occupancy code $O^{(t-1)}_{d+1}$ at $C^{(t)}_{d+1}+\delta$ by $e_{\delta}=E\!\big(O^{(t-1)}_{d+1}(C^{(t)}_{d+1}+\delta)\big)$. We average the neighbor descriptors over $\mathcal{N}_{d+1}$, and apply a linear projection $W_{t}$ to obtain a 32-dim feature:
@@
Table 1: Mean BPP (bits per point) over 300 frames (vox10). (– denotes not reported in the original paper.)
@@
Table 2: Computational complexity comparison on Loot & Redandblack (mean per frame). (– denotes not reported in the original paper.)
@@ 3.4 Entropy coding based on sibling correlation
Before encoding $G_{o}$, we summarize the already decoded even siblings into a lightweight context vector. Let $O_{e}(i)\in\{0,\ldots,255\}$ be the decoded 8-bit occupancy code for child $i\in G_{e}$. We embed each code with a 256-entry table $E\in\mathbb{R}^{256\times 32}$ and concatenate with its 3-dim relative position $\pi(i)\in\mathbb{R}^{3}$ (the child’s $(b_{x},b_{y},b_{z})$ inside the parent), then we apply a linear projection $W_{s}\in\mathbb{R}^{32\times(32+3)}$. The four descriptors are averaged with a mask (over available siblings) to form a sibling feature of the same width as $F_{d+1}$:
$F(G_{e})=\frac{1}{|G_{e}|}\sum_{i\in G_{e}}W_{s}\,[E(O_{e}(i)),\,\pi(i)],$ (9)
where $E(\cdot)$ is a 256-entry embedding table. For any child voxel $j\in G_{o}$, we fuse this context by element-wise addition for richer context $\tilde{F}_{d+1}$: $\tilde{F}_{d+1}=F_{d+1}+F(G_{e}).$ Then we reuse the same two prediction heads: first predict $s_{0}(j)$ from $\tilde{F}_{d+1}$, then embed $s_{0}(j)$, add it to $\tilde{F}_{d+1}$, and predict $s_{1}(j)$. Arithmetic coding is applied after each prediction. This keeps strict causality: odd children depend only on already decoded even siblings within the same parent.
@@ 4 Experiments
All experiments are run on a desktop with an Intel Core i9-14900 CPU and an NVIDIA RTX-4090 GPU. We implement the codec in PyTorch [10] with TorchSparse [15] for sparse 3D operations. We use the torchac [6] for entropy encoding.
@@
Main results Tab. 1 summarizes mean BPP at 10-bit precision on 8iVFBv2. HINT is competitive across all sequences: it achieves comparable compression ratio to ECM-OPCC and consistently outperforms G-PCC, SparsePCGC, OctAttention, ACNP-OctAttention, and VoxelDNN. Although with a slight rate gap to ECM-OPCC, Tab. 2 shows that HINT reduces encoding and decoding latency by one to two orders of magnitude compared with ECM-OPCC. This combination of compression and computational efficiency highlights HINT as a codec for practical deployment. Tab. 2 shows HINT runs at 105/140 ms (encode/decode) ($36.2\times$$29.3\times$ speedups over SparsePCGC and $49.6\times$/ $21.6\times$ over G-PCC), with a 3.4 MB (vs. 4.9 MB for SparsePCGC) model size.
@@
Ablation (Loot, mean over 300 frames). Removing temporal module increases BPP from 0.562 to 0.568, while removing sibling conditioning increases it to 0.587 (RENO: 0.592), suggesting sibling conditioning contributes more to the bitrate reduction. Meanwhile, temporal cues provide alignment-free gains, consistent with our low-latency design.